#3 Carbon tetrachloride it be heated with an elect
Answer #1

The rate of heat generation is 100 kW = 100 kJ / s.

So, heat generated in 1 hr is 100 kJ / s * 3600 s = 360000000 J.

75 kmol CCl4 contains 1000 * 75 mol = 75000 mol.

Specific heat capacity of CCl4 is 131.3 J/(mol K).

So, increase in temperature is 360000000 J / [75000 mol * 131.3 J/(mol K)] = 36.55 K = 36.55 oC.

So, the output temperature is (15 + 36.55) oC= 51.55 oC.